OpenStudy (erinweeks):

Find the vertical asymptotes, if any, of the graph of the rational function.

OpenStudy (erinweeks):

\[f(x) = \frac{ x }{ x ^{2} +1}\]

OpenStudy (erinweeks):

@Umangiasd could you help me with a few more?

OpenStudy (anonymous):

Sure, do you know limits?

OpenStudy (erinweeks):

A. x = 1 B. x = -1 C. x = -1, x = 1 D. no vertical asymptote

OpenStudy (erinweeks):

not really

OpenStudy (anonymous):

Okay, that will make it a little bit harder, the vertical asymptote exists when x indeterminate the function for a value of it As an example \[\frac{ 1 }{ x }\], when x=0 doesn't have a solution, because you can't divide by 0, so, there is a vertical asymptote of that function in yours \[f(x)= \frac{ x }{ x^2 +1 }\] so x^2 +1 = 0 for x=?

OpenStudy (erinweeks):

im guessing mine doesnt have one

OpenStudy (anonymous):

\[x^2+1=0 \leftarrow \rightarrow x^2=-1\] so, your function is continous for all R, and in fact, it hasn't a vertical asymptote c:

OpenStudy (erinweeks):

okay so it would be -1

OpenStudy (anonymous):

no, no, no x^2=-1 doesn't exist for real numbers, so, x^2+1 is never 0, that means you don't have a vertical asymptote xD